# Multiscale Analysis of the Conductivity in the Lorentz Mirrors Model

### v3

@@ 1 Introduction
The mirrors model, introduced in [8], provides a particularly striking instance of such a system. The scatterers are “mirrors” placed at the sites of $\mathbb{Z}^{d}$, each implementing a random reflection chosen uniformly among those satisfying reversibility and the no–U–turn constraint. The resulting dynamics is fully deterministic, time–reversible, and non–chaotic; in particular, infinitely many finite loops occur with positive probability. This leads to strong memory effects and non–Markovian behavior that severely complicate probabilistic analysis. In fact, even the question of whether trajectories are typically open or closed is unresolved in $d=3$ at full density.
Nevertheless, numerical experiments [3] suggest that a remarkable form of universality emerges at large scales: the model exhibits normal conductivity in $d=3$, with a crossing probability close to that of a non–backtracking random walk. This raises two natural questions: (i) how can a deterministic, non–chaotic system effectively behave as a Markov process at large scales? and (ii) can the conductivity be computed analytically?
@@ 5 Closure Assumption and Second-Order Terms
The correlation functions represent correlations between two crossings in opposite directions. Using (23), one can show (see Appendix C) that
@@ Appendix C Bounds on $R_{11}$, $R_{12}$, and $R_{2}$
Bound on $R_{2}$. Rewriting (22) and using translation invariance yields
$c_{n}(1-c_{n})^{2}R_{2}$ $=-2\sum_{y_{2}\in O_{1}^{-}}\mathbb{E}[\delta t_{n}(x_{1},y_{1})\delta t_{n}(x_{2},y_{2})]c_{n}(y_{1},x_{2}).$ (40)
Inserting the explicit form of correlations in the appropriate class $A_{cc}$ and discarding configurations where $c_{n}(y_{1},x_{2})=0$ by reversibility, we obtain
$c_{n}(1-c_{n})^{2}|R_{2}|\leq Ch(n)c_{n}^{2},$
for some constant $C>0$, hence
$|R_{2}|\leq\frac{Ch(n)}{2^{n}}.$
Together, these bounds imply the scaling form used in the main text for $\Delta_{n}$ and the resulting asymptotic recursion for $\kappa_{n}$.

### v4

@@ 1 Introduction
The mirrors model, introduced in [8], provides a particularly striking instance of such a system. The scatterers are “mirrors” placed at the sites of $\mathbb{Z}^{d}$, each implementing a random reflection chosen uniformly among those satisfying reversibility and the no–U–turn constraint. The resulting dynamics is fully deterministic, time–reversible, and non–chaotic; in particular, infinitely many finite loops occur with positive probability. This leads to strong memory effects and non–Markovian behavior that severely complicate probabilistic analysis. It is straightforward to see that $\bigl(\tfrac{p}{2d-1}\bigr)^{4}$ provides a lower bound on the probability that a particle starting from the origin becomes trapped in a finite loop forever. Proving true diffusivity of the motion in this model is therefore an elusive challenge. The best results in this direction are due to Elboim, Gloria, and Hernández [4], who showed that for all dimensions $d\geq 4$ and sufficiently small $p>0$, particle trajectories exhibit diffusive behavior up to times of order $1/p^{\alpha}$, with $\alpha>1$; in particular, trajectories remain open up to this time scale. In contrast, in dimension $d=2$ with $p=1$, it is known [6, 2] that every trajectory eventually closes with probability one, while the case $p<1$ was analyzed in [9]. In the present work we do not address diffusion or recurrence on $\mathbb{Z}^{d}$ itself but instead focus on establishing normal conductivity in the mirrors model. This property has been established in a dilute regime for a non-lattice random Lorentz gas [1], and in an anisotropic random lattice Lorentz gas at full density for $d\geq 7$ [7]. For this purpose, the most fundamental observable is the crossing probability of a finite slab [7, 3]. Despite the seemingly unfavorable features mentioned above from the viewpoint of diffusion, numerical simulations [3] for the case of full density ($p=1$) in $d=3$ support the conjecture that the mirrors model nevertheless displays normal conductive behavior. Specifically, the crossing probability $C_{N}$ of a slab of width $N$ appears to scale as $C_{N}\sim\kappa/N$, with $\kappa\simeq 1.535$, a value remarkably close to $3/2$. This coincides with the crossing probability of a non-backtracking random walk, suggesting that despite the strong memory effects inherent to the mirrors dynamics, its large-scale behavior remains close to that of a Markovian process. This raises two natural questions: (i) how can a deterministic, non–chaotic system effectively behave as a Markov process at large scales? and (ii) can the conductivity be computed analytically?
@@ 5 Closure Assumption and Second-Order Terms
Using (23), one can show (see Appendix C) that
@@ Appendix C Bounds on $R_{11}$, $R_{12}$, and $R_{2}$
Bound on $R_{2}$.We come now to $R_{2}$. By translation invariance, the sum over $x_{3}$ of the last factor produces a factor $c_{n}$ that may be taken out of the sum. Next the sum over $y_{2}$ that is running over elements of $O^{+}_{1}$ can be converted into a sum over $y_{2}\in O^{-}_{1}$ with a change of sign in front of the sum. This is because thanks to (7) we have the relations
$\sum_{y_{2}\in O^{+}_{1}}t_{n}(x_{2},y_{2})+\sum_{y_{2}\in O^{-}_{1}}t_{n}(x_{2},y_{2})=1$
$\sum_{y_{2}\in O^{+}_{1}}c_{n}(x_{2},y_{2})+\sum_{y_{2}\in O^{-}_{1}}c_{n}(x_{2},y_{2})=1$
The first relation is true for any realizations of the mirrors environment. Actually exactly one of the terms is equal to one, all the others are zero. The second line follows from the first one by taking expectation. So we are left with
$c_{n}(1-c_{n})^{2}R_{2}$
$=$ $-2\sum_{y_{2}\in O_{1}^{-}}\mathbb{E}[\delta t_{n}(x_{1},y_{1})\delta t_{n}(x_{2},y_{2})]c_{n}(y_{1},x_{2})$
Next, let us look at the sum above in more details
$\sum_{y_{2}\in O_{1}^{-}}\mathbb{E}[\delta t_{n}(x_{1},y_{1})\delta t_{n}(x_{2},y_{2})]c_{n}(y_{1},x_{2})$ (41)
$=$ $\sum_{y_{2}\in O_{1}^{-}}\left[\mathbb{E}[t_{n}(x_{1},y_{1})t_{n}(x_{2},y_{2})]-c_{n}(x_{1},y_{1})c_{n}(x_{2},y_{2})\right]c_{n}(y_{1},x_{2})$
Using then the positivity of the $c_{n}$ and the fact that $t_{n}\in\{0,1\}$, we have the easy upper bound $R_{2}\leq\frac{1}{1-c_{n}}$. Similarly a lower bound $R_{2}\geq-\frac{c_{n}}{1-c_{n}}$ is obtained so that
$|R_{2}|\leq\max\left(\frac{1}{1-c_{n}},\frac{c_{n}}{1-c_{n}}\right)$
Under the closure assumptions detailed above, we now derive a better bound on $R_{2}$. Observe first that the sums in (41) involves $(x_{1},y_{1},x_{2},y_{2})\in A_{cc}$. Therefore we plug (39) in (41). In this expression, we have $c_{n}(y_{1},x_{2})=\langle t_{n}(y_{1},x_{2})\rangle=0$ when $Ry_{1}=x_{2}$ by reversibility and bijectivity of the mirrors map. Thus,
$c_{n}(1-c_{n})^{2}R_{2}=-2\sum_{y_{2}\in O^{-}_{1}}h_{n,cc}(x_{1},y_{1},x_{2},y_{2})(1-\delta_{Ry_{2}x_{1}})c_{n}(x_{1},y_{1})c_{n}(x_{2},y_{2})$
and
$c_{n}(1-c_{n})^{2}|R_{2}|$ $\leq$ $4\sum_{y_{2}\in O^{-}_{1}}|h_{n,cc}(x_{1},y_{1},x_{2},y_{2})|c_{n}(x_{1},y_{1})c_{n}(x_{2},y_{2})$
$\leq$ $O(h(n))c_{n}^{2}$
and finally
$|R_{2}|\leq O(h(n))\frac{c_{n}}{(1-c_{n})^{2}}=\frac{O(h(n))}{2^{n}}$